The pH of a solution ofo.02 M (2-10 M) CH,oH (methanol or methyl alcoho) in pure water is: a. 1.70 b. 2.00 c. 2.30 d. 5.30 e, 7.00 f 8.70 g. 11.70 h. 12.30 i. None of the above, the correct pH value is
Answer #1

Ka of CH3OH = 3.2*10^-16

a)

for simplicity lets write weak acid CH3OH as HA

HA -----> H+ + A-

0.02 0 0

0.02-x x x

Ka = [H+][A-]/[HA]

Ka = x*x/(c-x)

since ka is small, x will be small and it can be ignored as compared to c

So, above expression becomes

Ka = x*x/(c)

so, x = sqrt (Ka*c)

x = sqrt ((3.2*10^-16)*0.02) = 2.53*10^-9

pH = -log [H+] = -log (2.53*10^-9) = 8.6

Answer: 8.60

you can go with option f